Todd Redden is a supervisor at Pinebank Construction. He contributes $2,688 annually to a FSA and has medical insurance that has a $250 monthly premium. He earns a salary of $66,400 and is paid semimonthly.


Required:
What is his taxable income per period? (Do not round intermediate calculations and round your final answer to 2 decimal places.)
Answer #1

Annual salary paid semimonthly $66400 / 24 = $2,766.67
• FSA contribution per pay period $2,688 / x 24 = $112.00
• Medical insurance per pay period $250 / 2 = $125
• $2,766.67 – 112 – 125 = $2,529.67

taxable income per period = $2529.67